OpenStudy (anonymous):

Solve using the square root method. 9x2 - 25 = 0 A. {-5/3, 5/3} B. {-3/5, 3/5} C. {-3, 5} D. {-5, 3} how would you solve this step by step

OpenStudy (anonymous):

9x^2 - 25 = 0 9x^2 = 25 (add 25 to both sides) 3x = ±5 (take the square root of both sides) x = 5/3 or x = -5/3 (divide by 3)

OpenStudy (anonymous):

\[x=\sqrt{25/9}\].....\[x=5/3\].....A

hero (hero):

They could have just said "solve it" and that would have been fine.